OpenStudy (agent47):

Please help!!!!! Ill fan and medal <3

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

......

OpenStudy (agent47):

Find the slope of the line perpendicular to y=mx-b+Int(e^x-x^2+3x-1, x, inf)

OpenStudy (anonymous):

Oh I cant do slope. sorry.

OpenStudy (agent47):

:( can anyone help me?

OpenStudy (agent47):

oops, typ-o y=mx-b+Int(e^x-x^2+3x-1, x, -inf, 0)

OpenStudy (shamil98):

lemme do it pls

OpenStudy (agent47):

i dont trust ur ansurs. Last one was wrong!

OpenStudy (shamil98):

LOL

OpenStudy (anonymous):

Let him do it, people make mistakes once in awhile, that is how we learn

OpenStudy (agent47):

okkkkk

OpenStudy (anonymous):

i don't even understand the question

OpenStudy (agent47):

Find the slope of the line perpendicular to \[y=mx-b+Integral(e^x-x^2+3x-1, \space x, -\infty, 0)\]

OpenStudy (anonymous):

\[y=mx+b+\int_{-\infty}^0(e^x-x^2+3x-1)dx\]

OpenStudy (agent47):

omg yes! sorry i couldnt get the integral sign to work ><

OpenStudy (anonymous):

if that is the question, then it is \(-\frac{1}{m}\) since the integral, if it converges, is a number

OpenStudy (anonymous):

here is the answer. −1/m i did it on a piece of paer.

OpenStudy (agent47):

how can i prove that the integral converges though?

OpenStudy (anonymous):

i don't think it converges however

OpenStudy (agent47):

satellite, can you fan me so that i can message you? I just need to send one message.

OpenStudy (anonymous):

maybe it is \[\large y=mx+b+\int_{-\infty}^0(e^{x-x^2+3x-1})dx\]